Draaksh Corporation sells premium quality wine for $120 per bottle. Its direct materials and direct labour costs are $23 and $13.00 respectively per bottle. It pays its direct labour employees a wage of $26 per hour. The company performed a regression analysis using the past 12 months’ data and established the following monthly cost equation for manufacturing overhead costs using direct labour hours as the overhead allocation base: y = $155,200 + $23.50x Draaksh believes that the above cost estimates will not substantially change for the next fiscal year. Given the stiff competition in the wine market, Draaksh budgeted an amount of $34,800 per month for sales promotions; additionally, it has decided to offer a sales commission of $6.25 per bottle to its sales personnel. Administrative expenses are expected to be $25,400 per month. Compute the expected total variable cost per bottle and the expected contribution margin ratio. 2. Compute the annual break-even sales in units and dollars. (Round your intermediate and final answers to the whole number.) 3. Draaksh has budgeted sales of $8.9 million for the next fiscal year. What is the company’s margin of safety in dollars and as a percentage of budgeted sales? (Round your intermediate and final answers to the whole number.)
Answer #1

Solution

Draaksh Corporation

Computation of expected total variable cost per bottle and the expected contribution margin ratio:

Expected total variable cost per bottle:

Direct materials

$23

Direct labor

$13

Variable manufautring overhead

$11.75

sales commission

$6.25

total variable cost per bottle

$54

contribution margin ratio:

contribution margin

sales price - variable cost

$120 - $54 = $66

Contribution margin ratio

$66/$120

55%

Note: variable manufacturing overhead is computed as follows,

Variable direct labor cost per unit = $13

Direct labor rate per hour = $26

Direct labor hours per unit = $13/$26 = 0.5 hours

The equation for MOH, Y= $155,200 + $23.50x

In the variable component of the equation, X = direct labor hours

Hence, variable overhead cost per unit = $23.50 x 0.5 hours = $11.75

  1. Annual break-even sales in units and dollars:

Break-even sales in units = fixed cost/contribution margin per unit

Fixed cost = MOH + administrative overhead + sales promotion

= 12 x ($155,200 + 25,400 + 34,800) = $2,584,800

Break-even sales in units = 2,584,800/66 = 39,164 units

Break-even point in sales dollars = fixed cost/contribution margin ratio

Annual Fixed cost = $2,584,800

Contribution margin ratio = 55%

Break-even sales in dollars = 2,584,800/55% = $4,699,636

  1. Computation of MOS –

Margin of safety = budgeted sales – break-even sales in dollars

Budgeted sales = $8,900,000

Break-even sales in dollars = $4,699,636

MOS = 8,900,000 – 4,699,636 = $4,200,364

MOS percentage –

Margin of safety percentage = (MOS sales/budgeted sales) x 100

= (4,200,364/8,900,000) x 100 = 47.2%